Random samples of size N=2 are drawn from a finite population consisting of the number 5,6,7,8, and 9. Compute for the mean, Variance,and Standard deviation, and also the mean, Variance, and the Standard deviation of the sample mean

1
Expert's answer
2022-04-11T10:05:03-0400

The population mean:

"\mu=\cfrac{5+6+7+8+9}{5}=7."


The population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\begin{Bmatrix}
 5-7,6-7,7-7, 8-7,9-7
\end{Bmatrix}="

"=\begin{Bmatrix}
-2, -1, 0, 1,2
\end{Bmatrix},"

"\sigma^2=(-2)^2\cdot \cfrac{1}{5}+(-1)^2\cdot \cfrac{1}{5}+0^2\cdot \cfrac{1}{5}+\\
+1^2\cdot \cfrac{1}{5}+2^2\cdot \cfrac{1}{5}=2."


The population standard deviation:

"\sigma=\sqrt{2}=1.414."


The mean of the sampling distribution of sample means:

"\mu_{\bar x} =\mu=7."


The variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{2}{2}=1."


The standard deviation of the sampling distribution of sample means:

"\sigma_{\bar x}=\cfrac{\sigma}{\sqrt n}=\cfrac{\sqrt 2}{\sqrt 2}=1."